19) If O.500 L of an unknown gas at STP has a mass of 1.70 g, what is its molar mass? Show detailed calculations
Answer #1

molar mass = (grams \times RT) / PV

STP = 1 atm of pressure and 273 K for temperature

R = 0.0821 atm L/mol K

molar mass = (1.7 g \times 0.0821 atm L/mol K \times 273 K ) / (1 atm \times 0.5 L)

molar mass = 76.20 g/mol